OpenStudy (anonymous):

What are the zeroes of the function? y = (x + 3)(x +2)(x – 2) A. –3, –2, 2 B. –3, –2, –2 C. 3, 2, 2 D. 3, 2, –2

OpenStudy (campbell_st):

to find the zeros all you need to do is solve x + 3 = 0 x + 2 = 0 x - 2 = 0 hope this helps